OpenStudy (anonymous):

Hey can anyone plz help me understand this question The percentage of salt in 1 l of water is 10%. If 500ml of water is added to this mixture what percentage of salt is there now. I am doing 10/100 times 1500

OpenStudy (xapproachesinfinity):

The added water is 1.5l not 1500 10/100 =1l X/100=1.5l X=15 percent

OpenStudy (anonymous):

No the answer is 6 times 2/3

OpenStudy (swamijatin2):

1l = 1000 ml ok if 500ml is added then 1000+500=1500 liters are there in 1000 liter % of salt is 10% then in 500 literes %of salt is half of 10 that is 5% so in 1500liters there is 15 % of salt

OpenStudy (xapproachesinfinity):

Does the added water have salt in it

OpenStudy (anonymous):

Hey @Orion1213 do u have any idea? Thnx for helping again

OpenStudy (anonymous):

initially with respect to volume of water: 1000 + 500 = 1500 considering % of salt solution: 0.1(1000) + 0 = 1500x x = 0.1(1000)/1500 = 0.0666666666667 in %: 100x = 6.67% = 6 and 2/3 percent....

OpenStudy (anonymous):

@Orion1213 how did u get 0.1(1000)

OpenStudy (anonymous):

10% of 1 li.(1000ml) of salt = 0.1(1000)

OpenStudy (anonymous):

Oh thank u

OpenStudy (anonymous):

np...\(\ddot\smile\)